Bobby is three years older than twice Arianna's age. If the total of their ages is 9, how old is Bobby, and how old is Arianna?
Answer by ankor@dixie-net.com(22740)

Bobby is three years older than twice Arianna's age.
b = 2a+3
If the total of their ages is 9,
a + b = 9
replace b with (2a+3)
a + 2a + 3 = 9
3a = 9 - 3
3a = 6
a = 6/3
a = 2 yrs is A's age
then
b = 2(2) + 3
b = 7 yrs is B's age